Quad Enterprises is considering a new 4-year expansion project that requires an initial fixed asset investment of $3.78 million. The fixed asset will be depreciated straight-line to zero over its 4-year tax life, after which time it will be worthless. The project is estimated to generate $3,360,000 in annual sales, with costs of $1,344,000.

  

If the tax rate is 21 percent, what is the OCF for this project?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(3,780,000/4)=$945000/year

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(3,360,000-1,344,000)(1-0.21)+(0.21*945000)

=$1,791,090.